A student scored 71, 83, and 96 on three algebra tests. What must he score on the fourth test in order to have an average of at least 85?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

A student scored 71, 83, and 96 on three algebra tests. What must he score on the fourth test in order to have an average of at least 85?
---
Let the 4th test score be "x":
Equation:
(71+83+96+x)/4 = 85
---
250+x = 4*85
x = 90
==========
Cheers,
Stan H.

Answer by solver91311(24713)



If you want the average of 4 numbers, you add the numbers and divide by 4. So if you have the average of 4 numbers, the sum of those numbers must be the average multiplied by 4.

Add the three scores already received. Multiply 85 by 4. Subtract the sum of the first three scores. The result will be the score required to achieve the stated average. You might want to compare the score required to the maximum score possible on the last test, if you know it. After all, it is possible that the stated average cannot be achieved.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism